# SDTrack: A Baseline for Event-based Tracking via Spiking Neural Networks

### v2

@@ 6 Conclusion
In this work, we propose the first transformer-based spike-driven tracking pipeline for event-based tracking. The pipeline consists of GTP method and SDTrack. GTP method provides stronger event representation by recording trajectory information, which we demonstrate to be robust and generalizable through extensive experimentation. SDTrack is the first transformer-based spike-driven tracker that learns enhanced positional information through IPL without requiring additional positional encoding. Notably, SDTrack is end-to-end and requires no data augmentation or post-processing, yet achieves performance comparable to state-of-the-art methods across datasets while maintaining minimal energy consumption and parameter counts. We hope this work serves as a solid baseline to inspire and facilitate future research.

### v3

@@ 6 Conclusion
In this work, we propose the first transformer-based spike-driven tracking pipeline for event-based tracking, consisting of the GTP method and SDTrack. GTP provides stronger event representation by recording trajectory information, demonstrating robustness and generalizability through extensive experimentation. SDTrack is the first transformer-based spike-driven tracker that learns enhanced positional information through IPL without requiring additional positional encoding. Notably, SDTrack is end-to-end and requires no data augmentation or post-processing, yet achieves performance comparable to state-of-the-art methods while maintaining minimal energy consumption and parameter counts. We hope this work serves as a solid baseline to inspire future research.